OpenStudy (anonymous):

To solve the equation m minus six equals thirteen, you should: subtract thirteen from both sides. subtract nineteen from both sides. add 6 to both sides. add thirteen to both sides.

OpenStudy (anonymous):

@MD152727

OpenStudy (md152727):

Hi

OpenStudy (md152727):

m-6=13 +6 +6 m=19 so C

OpenStudy (anonymous):

Solve. w minus four point eight equals five point one w equals nine point nine w equals five point five eight w equals 3 w equals zero point three

OpenStudy (md152727):

Please post in a new post as mediators say.